# Second-Order Fine-Tuning without Pain for LLMs: A Hessian Informed Zeroth-Order Optimizer

### v3

@@ Abstract
Fine-tuning large language models (LLMs) is necessary for specific downstream tasks, but the classic adaptive first-order optimizer entails prohibitive GPU memory because of backpropagation. Recent works such as MeZO have turned to zeroth-order optimizers for fine-tuning, which reduce substantial memory by using just two forward passes. However, heterogeneous curvatures across different parameter dimensions in LLMs often cause convergence instability or even failure. In this work, we propose HiZOO, a diagonal Hessian informed Zeroth-Order Optimizer , which is the first to leverage the diagonal Hessian to enhance ZOO for fine-tuning LLMs. We provide the theoretical proof for HiZOO and visualize the optimization trajectories on the test functions. Extensive experiments on various models (RoBERTa, OPT, Phi-2, and LLama3, with 350Msimilar-$\sim$66B parameters) indicate that HiZOO significantly reduces the number of training steps and improves model accuracy. For example, on the SST2 task, HiZOO achieves an 8x speed-up and better accuracy. Even when scaled to 66B-model, HiZOO outperforms MeZO with up to 5.1% absolute improvement. We also propose HiZOO-L, which reduces the Hessian memory cost to 10% of the MeZO, while maintaining almost same performance. Compared with ZO-Adam, HiZOO-L achieves a 4.3% absolute improvement, just using 50% of the GPU memory. Code is available at https://anonymous.4open.science/r/HiZOO-27F8.
Equal contribution. Correspondence to: yehaishan@xjtu.edu.cn
@@ 1 Introduction
Figure 1: (Left) Comparison of HiZOO, MeZO and Adam. (Right) The test example with heterogeneous curvatures. HiZOO updates along the direction with greater curvature ($X$) and converges more quickly than MeZO. The corresponding loss curves are shown in Section 3.5.
Fine-tuning pre-trained large language models (LLMs) for specific tasks has garnered significant attention. As the number of model parameter increases, full parameter fine-tuning (FT) becomes markedly memory-intensive. To alleviate the memory limitations, Parameter Efficient Fine-Tuning (PEFT) methods [19, 24, 10, 60, 35] have been developed, which only fine-tune a small number of (extra) model parameters. As a result, they significantly reduce computational and storage costs, while achieving performance comparable to a fully fine-tuned model.
First-order optimizers such as Adam [22] and AdamW [29] are widely used to fine-tune LLMs. Nevertheless, using these optimizers still leads to substantial memory consumption, primarily due to the inherent backpropagation process. To address these limitations, MeZO [32] proposed to utilize a zeroth-order optimizer (ZOO) to estimate the gradient with just two forward passes per step, no need for backpropagation anymore. This achieves numerous memory reduction and makes it accessible to train and store LLMs on consumer hardware.
However, parameters of LLMs often exhibit heterogeneous curvatures across different dimensions, as documented in recent studies [39, 16, 55]. This significant difference in curvature would make model converge towards the saddle point, slowing down the convergence speed, as shown in Figure 1 (right). Since the incorporation of Hessian to measure the curvature properties of the loss landscape, second order methods [26, 50, 2] can solve this sub-optimal behaviour. Unfortunately, in the context of zeroth-order optimization, one cannot directly compute the Hessian atop first-order derivatives.
In light of above, we propose HiZOO, as shown in Figure 1 (left), which estimates the diagonal Hessian by one more forward pass. HiZOO can act as a pre-conditioner, directly adjusting the update size of different parameters according to their curvatures. So that it can improve the model convergence when encountered with heterogeneous curvatures. As shown in Figure 2, HiZOO can significantly reduce the training steps and improve the accuracy. Here we summarize our key contributions as follows:
Figure 2: Performance of MeZO, HiZOO and HiZOO-L on SST2 task, when fine-tuning RoBERTa-large, OPT-13B, Llama3(8B) models. HiZOO can achieve 8$\times$ speedup and 1.55% absolute accuracy improvement compared with MeZO. Experiments are conducted with the same learning rate.
@@
    We conduct extensive experiments across different models (RoBERTa-large, OPT, Llama3 and Phi-2) with scales from 350M to 66B, different methods (FT, LoRA, prefix), and different downstream tasks (classification, multiple-choice, and generation) to verify the effect of the HiZOO. For example, on SST2 task HiZOO achieves a 1.55% accuracy improvement and 8×8$\times$8 × speedup over MeZO on average across different models. Even on OPT-66B, HiZOO outperforms better than MeZO with up to 5% improvement.
@@
    Further exploration in Section 4.3 showcases that HiZOO can achieve better performance in optimizing non-differentiable objectives such as F1 score. Specifically, HiZOO significantly outperforms the MeZO ’s results with 6.5% on average.
@@ 2 Related Works
First-Order Optimizer in LLMs    First-order optimizers, such as Gradient Descent (GD), Momentum, Adagrad [12], are fundamental in many areas like computer vision, natural languagle processing (NLP). Among them, Adam [22] plays a dominant role due to its fast convergence and is often chosen for training and fine-tuning LLMs. AdamW [29] improves upon Adam by adding the weight decay to alleviate overfitting. But both of them requires lots of memory cost due to the backpropagation process.
First-Order Optimizer Enhanced with Hessian    On the other hand, researchers incorporated second-order information (Hessian) to provide richer guidance for gradient descent during the training. For example BROYDEN [4] , Nesterov & Polyak [34] and Conn et al. [7] utilized curvature information to pre-condition the gradient; Magoulas et al. [31] applied diagonal Hessian as the pre-conditioner; Martens [33] approximated the Hessian with conjugate gradient. Sophia [26] used a light-weight estimate of the diagonal Hessian for pre-training LLMs. Despite their potential, above optimizers require the enormous GPU-memory cost.
Zeroth-Order Optimizer    Zeroth-order optimizers, with just forward passes, can greatly reduce the memory consumption. Methods like SPSA [42] have been shown to perform well in non-convex multi-agent optimization [44, 18] or generating black-box adversarial examples [6, 5, 27, 52]. Recently, MeZO [32] firstly adapted the classical ZO-SGD method to fine-tune LLMs, achieving comparable performance with significant memory reduction. But these methods often struggle with heterogeneous curvatures.
@@ 3 Methods
In the following, we briefly introduce the classical ZO gradient estimator SPSA [42], which is used in MeZO. Then we describe how HiZOO estimates diagonal Hessian and cooperates with ZOO .
@@ 3.1 Preliminaries
(Simultaneous Perturbation Stochastic Approximation or SPSA). Given a model with parameters $\theta\in\mathbb{R}^{d}$ and loss function $\mathcal{L}$, SPSA estimates the gradient on a minibatch $\mathcal{B}$ as:
@@ 3.2 Hessian Informed Zeroth-Order Optimization
19:
20:function PerturbParameter($\theta$, $\mu$, $\Sigma^{1/2}_{t}$, $s$)
21:     Reset random number generator with seed $s$triangleright For sampling $u_{i}$
22:     for $\theta_{i}\in\theta$ do
23:         Sample $u_{i}\sim\mathcal{N}(0,I_{d})$
24:         $\theta_{i}\leftarrow\theta_{i}+\mu\Sigma^{1/2}_{t}u_{i}$triangleright Modify parameters in place
25:     end for
26:     return $\theta$
27:end function
@@ 3.3 Diagonal Hessian Estimator
Storing the exact full spectral Hessian ($d\times d$) requires $\mathcal{O}(d^{2})$ memory [49, 48, 9], which is sufficient but never necessary. In HiZOO, we just estimate and retain only the diagonal Hessian which requires $\mathcal{O}(d)$ memory. It serves as a pre-conditioner to scale the direction and magnitude of the model parameter updates according to their respective curvatures.
Drawing from the lemma presented in MiNES [51]:
@@
where $H$ is the Hessian $\nabla^{2}\mathcal{L}(\theta;\mathcal{B})$ and $\Sigma$ is a positive definite matrix with $\Sigma^{-1}\approx\nabla^{2}\mathcal{L}(\theta;\mathcal{B})$.
@@
$\mathcal{L}(\theta+\mu\Sigma^{1/2}u;\mathcal{B})=\mathcal{L}(\theta;\mathcal{B})+\mu\langle\mathcal{L}(\theta;\mathcal{B}),\Sigma^{1/2}u\rangle+\frac{\mu^{2}}{2}u^{\top}\Sigma^{1/2}\nabla^{2}\mathcal{L}(\theta;\mathcal{B})\Sigma^{1/2}u+\alpha(\theta,\mu\Sigma^{1/2}u).$ (4)
@@
Upon substituting the above results into the left side of the Eq. (3), we arrive at:
@@
To further reduce Hessian memory consumption, we propose HiZOO-L to maintain it in a low-rank subspace, motivated by Adafactor [41]. For $\hat{\Sigma}^{-1}\in\mathbb{R}^{p\times q}$, we will store two low-rank matrices $R\in\mathbb{R}^{p\times k}$ and $C\in\mathbb{R}^{k\times q}$ with $k=1$. Specifically, we can get $\hat{\Sigma}^{-1}$ by:
@@ 3.4 Convergence Analysis
$g_{\mu}(\theta_{t})=\sum_{i=1}^{b}\frac{\mathcal{L}(\theta_{t}+\mu\Sigma_{t}^{1/2}u_{i};\mathcal{B})-\mathcal{L}(\theta_{t}-\mu\Sigma_{t}^{1/2}u_{i};\mathcal{B})}{2b\mu}\Sigma_{t}^{1/2}u_{i}.$ (7)
@@
$\mathbb{E}\left[\mathcal{L}(\theta_{t+1};\mathcal{B})\mid\right]\leq\mathcal{L}(\theta_{t};\mathcal{B})-\frac{\eta_{t}}{4}\|\nabla\mathcal{L}(\theta_{t};\mathcal{B})\|_{\Sigma_{t}}^{2}+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}+\mathcal{O}(\mu^{2}).$ (8)
@@
$\begin{split}\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{\mbox{out}};\mathcal{B})\|^{2}\right]\leq\frac{32L\left(\max_{t}\{\mathrm{tr}(\Sigma_{t})\}+\beta_{u}\right)(\mathcal{L}(\theta_{1};\mathcal{B})-\mathcal{L}(\theta_{*};\mathcal{B}))}{\sqrt{T}\beta_{\ell}}+\frac{\sigma^{2}}{T^{3/2}\beta_{\ell}}+\mathcal{O}\left(\mu^{2}\right),\end{split}$ (9)
where $\mathcal{L}(\theta_{*};\mathcal{B})$ minimizes the function $\mathcal{L}(\theta;\mathcal{B})$. The above equation shows that as $T\to\infty$, HiZOO can converge to the optimum point.
@@ 4 Experiments
Large language models are generally classified into two types: (1) Encoder-Decoder, also known as masked language models, such as BERT [11] and ALBERT [23]; (2) Decoder-Only, also recognized as generative language models, such as GPT family [37, 3], OPT [57], LLaMA [45], Phi [25, 17].
To rigorously assess the universality and robustness of our HiZOO, we have chosen models from each category for empirical testing. Additionally, we investigate FT and PEFT (LoRA [19] and prefix [24]). Detailed experiment settings are presented in Appendix E.1.
@@ 4.1 Masked Language Models
Firstly, we conduct experiments on RoBERTa-large 350M [28] on three NLP task paradigms: sentence classification, multiple choice and text generation. We follow the experimental setting [32] in studying the few-shot and many-shot, sampling $k$ examples per class for $k=16$ (results in Table 1), which can reach the following observations and summaries.
@@
Table 1: Experiments on RoBERTa-large (350M parameters, k=16). PEFT represents the LoRA and prefix and we report the best of them. All reported numbers are averaged accuracy (standard deviation) across 5 runs.
Figure 4: OPT-13B results with zero-shot, in-context learning (ICL), MeZO, HiZOO and fine-tuning with Adam (FT). PEFT represents the best among LoRA and prefix. More details can be found in Appendix E.2.
HiZOO greatly increases the convergence speed across full-parameter tuning, LoRA and prefix. As shown in Figure 4, HiZOO achieves 4×4$\times$4 × speedup over MeZO on average while getting the same training loss compared with MeZO. What’s more, HiZOO finally achieves a 1.57% accuracy improvement better than MeZO. More loss curves can be seen in Appendix E.1.
HiZOO achieves better performance compared with MeZO. Table 1 shows that HiZOO outperforms MeZO’s results with 3.5% on average on all datasets across different tasks. Specifically, HiZOO outperforms MeZO more than 6% in accuon SNLI and MNLI dataset.
4.2 Auto-Regressive Language Models
Then we extend experiments with OPT family (13B, 30B, 66B) on the same NLP task paradigms. The main results on OPT-13B are presented in Table 2. We can see that HiZOO boosts MeZO’s accuracy for 1.1% accuracy on average using prefix-tuning. Specifically, HiZOO achieves 2.87% and 3% accuracy improvement on WSC and COPA tasks.
HiZOO is capable of scaling to large models with up to 66B parameters, while preserving its exceptional performance. As depicted in Table 4, on OPT-30B HiZOO outperforms MeZO with up to 2.9% increase and 0.8% increase on average. Even scaling to OPT-66B, HiZOO(prefix) still outperforms MeZO(prefix) with up to 5.1% increase and 2% increase on average.
In addition, We utilize HiZOO to fine-tune the newly proposed model, Phi-2 (2.7B) and Llama3 (8B). Experiment results in Table 2 show that HiZOO outperforms MeZO in most cases. We also provide relative loss curves to show the better convergence process of our HiZOO in Appendix E.3.

### v4

# Equal contribution. Correspondence author

@@ Abstract
Fine-tuning large language models (LLMs) is necessary for specific downstream tasks, but the classic adaptive first-order optimizer entails prohibitive GPU memory because of backpropagation. Recent works such as MeZO have turned to zeroth-order optimizers for fine-tuning, which reduce substantial memory by using just two forward passes. However, heterogeneous curvatures across different parameter dimensions in LLMs often cause convergence instability or even failure. In this work, we propose HiZOO, a diagonal Hessian informed Zeroth-Order Optimizer , which is the first to leverage the diagonal Hessian to enhance ZOO for fine-tuning LLMs. We provide the theoretical proof for HiZOO and visualize the optimization trajectories on the test functions. Extensive experiments on various models (RoBERTa, OPT, Phi-2, and LLama3, with 350Msimilar-$\sim$66B parameters) indicate that HiZOO significantly reduces the number of training steps and improves model accuracy. For example, on the SST2 task, HiZOO achieves an 8$\times$ speed-up and better accuracy. Even when scaled to 66B-model, HiZOO outperforms MeZO with up to 5.1% absolute improvement. We also propose HiZOO-L, which reduces the Hessian memory cost to 10% of the MeZO, while maintaining almost same performance. Compared with ZO-Adam, HiZOO-L achieves a 4.3% absolute improvement, just using 50% of the GPU memory. Code is available at https://github.com/Yanjun-Zhao/HiZOO.
Figure 1: (Left) Comparison of HiZOO, MeZO and Adam. (Right) Heterogeneous curvatures example. HiZOO updates along the direction with greater curvature ($X$) and converges more quickly than MeZO. The corresponding loss curves are shown in Section 3.5.
@@ 1 Introduction
Fine-tuning pre-trained LLMs for specific tasks has gained significant attention recently. As the number of model parameters increases, full parameter fine-tuning (FT) becomes markedly memory-intensive. To alleviate GPU memory limitations, parameter-efficient fine-tuning (PEFT) methods (Hu et al., 2022; Li & Liang, 2021; Dettmers et al., 2023; Zhao et al., 2024b; Pan et al., 2024) have been developed, which only fine-tune a small number of (extra) model parameters. As a result, they significantly reduce the computational and storage cost, while achieving performance comparable to a fully fine-tuned model.
Adaptive first-order optimizers such as Adam (Kingma & Ba, 2015) and AdamW (Loshchilov & Hutter, 2019) are widely used to fine-tune LLMs. However, using these optimizers still leads to substantial memory consumption, primarily due to the inherent backpropagation process to calculate the gradient. To address these limitations, MeZO (Malladi et al., 2023) proposed to utilize a zeroth-order optimizer (ZOO) to estimate the gradient with just two forward passes per step, no need for backpropagation anymore. This achieves numerous memory reductions and makes it accessible to train and store LLMs on consumer hardware.
However, the parameters of LLMs often exhibit heterogeneous curvatures across different dimensions, as documented in recent studies (Sagun et al., 2017; Ghorbani et al., 2019; Zhang et al., 2020). This significant difference of second derivative makes the MeZO converge towards saddle point, slowing down the convergence speed, as shown in Figure 1 (right). Since the incorporation of Hessian to measure the curvature properties of the loss landscape, second-order methods (Liu & Li, 2023; Yao et al., 2021; Anil et al., 2021) can solve this suboptimal behavior. Unfortunately, in the context of zeroth-order optimization, one cannot directly compute the Hessian atop first-order derivatives.
Figure 2: Performance of MeZO, HiZOO and HiZOO-L on SST2 task, when fine-tuning RoBERTa-large, OPT-13B, Llama3(8B) models. HiZOO can achieve 8$\times$ speedup and 1.55% absolute accuracy improvement compared with MeZO.
In light of above, we propose HiZOO, as shown in Figure 1 (left), which estimates the diagonal Hessian by one more forward pass. HiZOO can act as a pre-conditioner, directly adjusting the update size of different parameters according to their curvatures. So that it can improve the model convergence when encountered with heterogeneous curvatures. As shown in Figure 2, HiZOO can significantly reduce number of training steps and improve model accuracy. Here we summarize our key contributions as follows:
@@
    We conduct extensive experiments across different models (RoBERTa-large, OPT, Llama3 and Phi-2) with scales from 350M to 66B, different methods (FT, LoRA, prefix), and different downstream tasks (classification, multiple-choice, and generation) to verify the effect of the HiZOO. For example, on SST2 task HiZOO achieves a better accuracy and 8$\times$ speedup over MeZO on average across different models. Even on OPT-66B, HiZOO outperforms better than MeZO with up to 5.1% absolute improvement.
@@
    Further exploration in Section 4.3 showcases that HiZOO can achieve better performance in optimizing non-differentiable objectives such as F1 score. Specifically, HiZOO significantly outperforms MeZO ’s results with 6.5% absolute on average.
@@ 2 Related Works
First-Order adaptive optimizer used in fine-tuning LLMs    Optimization methods have consistently been a popular research domain. Adaptive first-order optimizer, such as Gradient Descent (GD), Momentum, Adagrad (Duchi et al., 2011), are fundamental in many areas like computer vision, natural languagle processing (NLP). Among them, Adam (Kingma & Ba, 2015) plays a dominant role due to its fast convergence and is often chosen for training and fine-tuning LLMs. AdamW (Loshchilov & Hutter, 2019) improves upon Adam by adding the weight decay to alleviate overfitting. But both of them requires lots of memory cost due to the backpropagation process. This issue has become increasingly critical as the number of LLM parameters skyrockets.
Enhanced optimizers with Hessian    On the other hand, researchers incorporated second-order information (Hessian) to provide richer guidance for gradient descent during the training. For example BROYDEN (BROYDEN, 1970) , Nesterov & Polyak (Nesterov & Polyak, 2006) and Conn et al. (Conn et al., 2000) utilized curvature information to pre-condition the gradient; Magoulas et al. (Magoulas et al., 1999) applied diagonal Hessian as the pre-conditioner; Martens (Martens, 2010) approximated the Hessian with conjugate gradient. Sophia (Liu & Li, 2023) used a light-weight estimate of the diagonal Hessian for pre-training LLMs. Despite their potential, above optimizers require the enormous GPU-memory cost. Additionally, these methods can only be used when first-order gradients are available.
Zeroth-Order Optimizer    Zeroth-order optimizers, with just forward passes to estimate the gradient, can greatly reduce the memory consumption. It appears in a wide range of applications where either the objective functions is implicit or its gradient is impossible or expensive to obtain. Methods like SPSA (Spall, 1992) have been shown to perform well in non-convex multi-agent optimization (Tang et al., 2021; Hajinezhad & Zavlanos, 2018) or generating black-box adversarial examples (Chen et al., 2017; Cai et al., 2021; Liu et al., 2019a; Ye et al., 2019). Recently, MeZO (Malladi et al., 2023) first adapted the classical ZO-SGD method to fine-tune LLM, achieving comparable performance with significant memory reduction. Then Zhang et al. (2024) proposed a wider array of ZO optimization techniques. However, these methods often struggle with heterogeneous curvatures.
@@ 3 Methods
In the following, we briefly introduce the classical ZO gradient estimator SPSA (Spall, 1992), which is used in MeZO. Then we describe how HiZOO estimates diagonal Hessian and cooperates with ZOO. We also provide detailed proof for our method.
@@ 3.1 Preliminaries
Simultaneous Perturbation Stochastic Approximation or SPSA
Given a model with parameters $\theta\in\mathbb{R}^{d}$ and loss function $\mathcal{L}$, SPSA estimates the gradient on a minibatch $\mathcal{B}$ , based on the concepts of sampling and differencing, as shown below:
@@ 3.2 Hessian Informed Zeroth-Order Optimization
19:function PerturbParameter($\theta$, $\mu$, $\Sigma^{1/2}_{t}$, $s$)
20:     Reset random number generator with seed $s$triangleright For sampling $u_{i}$
21:     for $\theta_{i}\in\theta$ do
22:         Sample $u_{i}\sim\mathcal{N}(0,I_{d})$
23:         $\theta_{i}\leftarrow\theta_{i}+\mu\Sigma^{1/2}_{t}u_{i}$triangleright Modify parameters in place
24:     end for
25:     return $\theta$
26:end function
@@ 3.3 Diagonal Hessian Estimator
Given a model with parameters $\theta\in\mathbb{R}^{d}$, storing the exact full spectral Hessian ($d\times d$) requires $\mathcal{O}(d^{2})$ memory (Yao et al., 2018; Xu et al., 2019; Dembo et al., 1982), which is sufficient but never necessary. In HiZOO, we just estimate and retain only the diagonal Hessian which requires $\mathcal{O}(d)$ memory. It serves as a pre-conditioner to scale the direction and magnitude of the model parameter updates according to their respective curvatures.
Drawing from the lemma presented in MiNES (Ye, 2023):
@@
where $H$ is the Hessian $\nabla^{2}\mathcal{L}(\theta;\mathcal{B})$ and $\Sigma$ is a positive definite matrix.
@@
$\mathcal{L}(\theta+\mu\Sigma^{1/2}u;\mathcal{B})=\mathcal{L}(\theta;\mathcal{B})+\mu\langle\mathcal{L}(\theta;\mathcal{B}),\Sigma^{1/2}u\rangle+\frac{\mu^{2}}{2}u^{\top}\Sigma^{1/2}\nabla^{2}\mathcal{L}(\theta;\mathcal{B})\Sigma^{1/2}u+\alpha(\theta,\mu\Sigma^{1/2}u).$
@@
Upon substituting the above results into the left side of the Eq. equation 3, we arrive at:
@@
To further reduce Hessian memory consumption, we propose HiZOO-L to maintain it in a low-rank subspace, motivated by Adafactor (Shazeer & Stern, 2018). For $\hat{\Sigma}^{-1}\in\mathbb{R}^{p\times q}$, we will store two low-rank matrices $R\in\mathbb{R}^{p\times k}$ and $C\in\mathbb{R}^{k\times q}$ with $k=1$. Specifically, we can get $\hat{\Sigma}^{-1}$ by:
@@ 3.4 Convergence Analysis
$g_{\mu}(\theta_{t})=\sum_{i=1}^{b}\frac{{\bm{L}}(\theta_{t}+\mu\Sigma_{t}^{1/2}u_{i};\mathcal{B}_{t})-{\bm{L}}(\theta_{t}-\mu\Sigma_{t}^{1/2}u_{i};\mathcal{B}_{t})}{2b\mu}\Sigma_{t}^{1/2}u_{i}.$ (6)
@@
$\mathbb{E}\left[{\bm{L}}(\theta_{t+1})\mid\right]\leq{\bm{L}}(\theta_{t})-\frac{\eta_{t}}{4}\|\nabla{\bm{L}}(\theta_{t})\|_{\Sigma_{t}}^{2}+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}+\mathcal{O}(\mu^{2}).$ (7)
@@
$\begin{split}\mathbb{E}\left[\|\nabla{\bm{L}}(\theta_{\mbox{out}})\|^{2}\right]\leq\frac{32L\left(\max_{t}\{\mathrm{tr}(\Sigma_{t})\}+\beta_{u}\right)({\bm{L}}(\theta_{1})-{\bm{L}}(\theta_{*}))}{\sqrt{T}\beta_{\ell}}+\frac{\sigma^{2}}{T^{3/2}\beta_{\ell}}+\mathcal{O}\left(\mu^{2}\right),\end{split}$ (8)
where ${\bm{L}}(\theta_{*})$ minimizes the function ${\bm{L}}(\theta;)$. The above equation shows that as $T\to\infty$, HiZOO can converge to the stationary point.
@@ 4 Experiments
Large language models are generally classified into two types: (1) Encoder-Decoder, also known as masked language models, such as BERT (Devlin et al., 2019) and ALBERT (Lan et al., 2020); (2) Decoder-Only, also recognized as generative language models, such as GPT family (Radford et al., 2019; Brown et al., 2020), OPT (Zhang et al., 2022a), LLaMA (Touvron et al., 2023), Phi (Li et al., 2023; Gunasekar et al., 2023).
To rigorously assess the universality and robustness of our HiZOO, we have chosen models from each category for empirical testing. Additionally, we investigate FT and PEFT (LoRA (Hu et al., 2022) and prefix (Li & Liang, 2021)). Detailed experiment settings are presented in Appendix E.1.
@@ 4.1 Masked Language Models
Firstly, we conduct experiments on RoBERTa-large 350M (Liu et al., 2019b) on three NLP task paradigms: sentence classification, multiple choice and text generation. We follow the experimental setting (Malladi et al., 2023) in studying the few-shot and many-shot, sampling $k$ examples per class for $k=16$ (results in Table 1) and $k=512$ (results in Appendix E.1). We did not utilize HiZOO-L here due to model’s smaller parameter count.
Table 1: Experiments on RoBERTa-large (350M parameters, k=16). PEFT represents using LoRA and prefix and we report the best result of them. All reported numbers are averaged accuracy (standard deviation) across 5 runs.
@@
Figure 4: Training loss curves when using Adam, MeZO and HiZOO to fine-tune Roberta-large on MNLI. The evaluation accuracy curves can be found in Figure 11 in Appendix E.1.
HiZOO greatly increases the convergence speed across full-parameter tuning, LoRA and prefix. As shown in Figure 4, HiZOO achieves 4$\times$ speedup over MeZO on average while getting the same training loss compared with MeZO. What’s more, HiZOO finally achieves a 2.2% absolute accuracy improvement on MNLI better than MeZO.
HiZOO achieves better performance compared with MeZO. Table 1 shows that HiZOO outperforms MeZO’s results with 3.5% absolute on average on all datasets across different tasks. Specifically, HiZOO outperforms MeZO more than 6% in both the SNLI and MNLI dataset.
@@
4.2 Auto-Regressive Language Models
Then we extend experiments with Phi-2(2.7B), Llama3(8B) and OPT family on the same NLP task paradigms. The results of the experiment in Table 2 show that HiZOO outperforms MeZO in most cases. Also, we can see that HiZOO-L has only a slight decrease in accuracy. We also provide relative loss curves to show the better convergence process of our HiZOO in Appendix E.2.
HiZOO is capable of scaling to large models with up to 66B parameters, while preserving its exceptional performance. As depicted in Table 4, on OPT-30B HiZOO outperforms MeZO with up to 2.9% increase and 1.1% increase on average. Even scaling to OPT-66B, HiZOO(prefix) still outperforms MeZO(prefix) with up to 5.1% increase and 2.7% increase on average.